### v1

# Emergence of Fluctuation Relations in UNO

Abstract
Fluctuation theorems are generalisations of the second law that describe the relations between work, temperature, and free energy in thermodynamic processes and are used extensively in studies of irreversibility and entropy. Many experiments have verified these relations for different physical systems in the setting of thermodynamics. In this study, we observe the same behavior away from physical thermodynamics, namely for the card game UNO, by performing numerical simulations of the game. As the analog of work, we choose the number of steps one player needs to effect a transition in her deck; the other players and the remaining cards play the role of a finite, non-Markovian bath. We also compare our observation with is expected for a Markovian random walk.
I Introduction
Fluctuation relations have become a mainstay in a variety of fields in science. They are associated most with stochastic thermodynamics [1, 2, 3, 4, 5, 6, 7], but also feature in fields like biophysics, optomechanics and quantum thermodynamics [8, 9, 10, 11, 12, 13]. Their significance lies in their generalization of the second law, being applicable even when systems are driven far out-of-equilibrium under non-quasistatic transformations. They describe how, despite a lack of thermodynamic constraints, strong relationships connecting average work, temperature and free energy hold. For this reason, fluctuation relations are invoked as measures of irreversibility and entropy, beyond the thermodynamic regime [12, 14, 15].
It has also become widely accepted that much of the underlying structure of fluctuation relations comes from logical and mathematical necessity [16, 17, 18]. The physical features of the system and whether some scenario adheres to some specific rendition of a fluctuation relation (say, Crooks’ theorem) boil down to assumptions. These are expressed in terms of detailed balance, or, quite equivalently, the encoding of a reference prior that connects the forward and reverse processes [5, 17, 18, 19].
Now, with this in mind, the following question emerges: if it is the case that fluctuation relations come from logical consistency, with the physics embedded in priors and assumptions – then it seems reasonable to expect some non-physics scenarios to obey certain fluctuation relations. There are plenty of possible contexts that may answer this question. For this report, we focus on games. In particular, an extensive-form scenario with moves-by-nature that is expressed as a card-shedding game [20, 21]. We are talking about the popular card game, UNO. We plot fluctuation relation graphs for various transitions in hand-sizes, in the context of this game. In particular, we found that when a specific class of cards are omitted from the game, most transitions adhere to Crooks’ fluctuation theorem.
We section the report in the following way:
  • In Section II, we review the rules for UNO. For those who are familiar with how the game works can skip this section entirely.
  • In Section III, we do two things. First, we run through the conceptual building blocks of a fluctuation relation experiment. Secondly, we show how we translate this to the game-theoretic context of UNO. We also clarify how our results are obtained in a non-trivial way.
  • In Section IV, we include the key results to be discussed later, noting standout features that should be explained.
  • In Section V, we provide explanations for the features raised in the previous section. In particular, we recontextualize thermodynamic notions (like temperature) to the game-theoretic setting.
  • Finally, in Section VI, we consolidate our key takeaways and conclude.
II Reviewing the Rules of UNO
Figure 1: The setup of a typical UNO game.
UNO (from Spanish and Italian for ‘one’), is a proprietary American card game developed in 1971 by Merle Robbins and is currently owned by Mattel, Inc. It is a shedding-type card game where players take turns to play a card onto a game pile, with the objective of losing all the cards in one’s hand. It is a derivative of a game from the 1940s known as Eights or Crazy Eights  [^1] These games are likewise a derivative of an even older card-shedding game called Mau-Mau.. These two earlier games could be played with a standard poker deck (sometimes with certain cards removed). UNO is played with a specially printed deck.
There are many variations to the game and house rules are very common. In this paper, we will go with the official rules [23].
The game consists of the deck (face-down), the gamepile (face-up), and the players’ hands (known only to the respective owner). During a player’s turn, they must choose and play a card from their hand onto the game pile, according to the card on top of the game pile. The card played must match either the same colour, or the same number/symbol as this topmost card. If this is not possible and there are no other legal moves, the player must draw one (and only one) card from the deck.
The game starts with every player being dealt a hand of seven cards, drawn randomly from the deck. The topmost card from the remaining deck is drawn onto the game pile. Once the starting player is designated, the turn order begins clockwise  [^2] There are official rules about how the dealer and the first player is chosen and so on. There are also rules about what to do when your handsize is reduced to one. These are irrelevant for this work.. The first player to lose all their cards wins.
We move on to the deck composition. The standard UNO deck consists of 108 cards:
  • Normal Cards: These cards are separated into four colours (blue, green, red, and yellow). Each card has a number between 0 to 9 and there are two copies of each, except for 0 which only has one, for a total of seventy-six cards.
  • Special Cards: These cards come in the same four colours but they enforce special actions or rulings when played. These come in three types:
    • The skip card skips the next player.
    • The reverse card reverses the direction of the play, i.e. it inverts the turn order.
    • The +2 card forces the next player to draw two cards from the deck and skip their turn.
    There are two copies of each, for a total of twenty-four cards.
  • Wildcards: These cards do not feature any specific colour or number. Rather, they allow the player to decide the next colour to be matched. The first variation of wildcard does just this, while a second variation also invokes the ruling of the +2 card, except it adds four cards instead of two. During a player’s turn, they can always play a wildcard in their hand regardless of what the game pile’s topmost card is. There are four copies of each variation, for a total of eight cards.
Of special note are the +2 cards and the +4 wildcards. We shall refer to them as plus-cards in this work. They will be shown to affect how much the game adheres to Crooks’ fluctuation relation, which we will now introduce.
III Methods
III.1 Key Elements of a Fluctuation Relations Experiment
We shall derive a detailed fluctuation theorem. The first such example was found by Evans and Searles [25]. Shortly later, Crooks derived his famous one, that reads [3]
$\frac{P_{f}(W)}{P_{r}(-W)}=e^{\beta(W-\Delta F)}.$ (1)
On the l.h.s. is the ratio between the distribution of work $W$ observed in the forward process and its associated reverse process, while the r.h.s. relates to the physics of the process. Crooks derived it for a purely deterministic (Hamiltonian) drive between two Gibbs state with inverse temperature $\beta$ and difference in free energy $\Delta F$. The same expression was then derived for the isothermal drive (Hamiltonian drive while the system exchanges energy with a thermal bath). This became the standard introductory scenario, see e.g. [5]. But detailed fluctuation theorems appear in many varieties: for instance, by conditioning on an observed value of some other parameter [26]; or independently for separate contributions to the total entropy production [27]; and many more. The version of Crooks’ theorem that we shall use is
$\frac{P_{f}(W|x\rightarrow y)}{P_{r}(-W|y\rightarrow x)}=e^{\beta(W-W_{0})}$ (2)
where $x$ is the initial (final) state of the target system for the forward (reverse) process, $y$ is the final (initial) state; $W$ designates work done on the target; $W_{0}$ is an offset for the origin of work ($\Delta F$ in thermal setting, but we won’t have this interpretation here).
In fact, any variable (typically entropy production), which can be written as the log-ratio between the statistics of the forward and the reverse process, obeys automatically the detailed fluctuation theorem [28, 17, 29]. To avoid falling into this triviality, experimental tests of detailed fluctuation theorem must define independently the variable to be measured (work) and the recipe for the reverse process. In the paradigmatic biophysical setting [8], the forward process being pulling a folded RNA strand till is straightens, the reverse process consists in pushing it; meanwhile, the work made on the system is measured as the energy invested to make the transition happen. We shall follow similar prescriptions in this paper.
III.2 Getting Fluctuation Relations for UNO
Figure 2: In the forward process, Alice follows a protocol that is effective at shedding cards. Meanwhile, in the reverse process Alice does the opposite of whatever the forward protocol entails. See Section III.2.4 and Appendix A for details.
(a) UNO under official rules
(b) UNO under official rules, but with +2 cards and +4 variation wildcards removed.
Figure 3: Two sets of Crooks graphs for settings stated above. The data used to plot the graphs consist of $4\times 10^{6}$ UNO games, split evenly between the two different versions of the game and the forward and reverse processes in each of them. The lines are the Theil–Sen regression lines for each of the transitions, which is more robust to outliers than the simple linear regression [30, 31]. Data points which have a sample size of smaller than 10 have been removed. We note the oscillatory behaviour at small values of $W$.
Having reviewed the key elements of fluctuation theorem experiments, we move to translating them into our UNO setting.
III.2.1 Target System & Bath
For studies of fluctuation theorems, the target system and the heat bath need to be identified. This distinguishes the forward and reverse processes, demarcating what deterministic work is being done on and where stochastic heat is coming from. For the experiment in [8], the target and the bath would be the RNA strand and the wet environment respectively.
For this work’s UNO games, we can isolate one player, labelled as Alice, as the target system. The game pile, main deck and the other players qualify as a stochastic context, acting like a bath contributing randomness into Alice’s states. It is a finite bath with memory, i.e. non-Markovian, because the deck is finite and (anti-) correlated with the system (the cards comprising the bath are those that are not with Alice).
In our simulations, all the other players in the bath are playing randomly, as opposed to Alice, which always play with a strategy (see Sec. III.2.4). The impact of changing the behaviour of the other players can be seen in Appendix D.
III.2.2 Initial & Final States
Now, what are “Alice’s states”? For our context, we choose hand-sizes (that is, how many cards are in Alice’s hand) as what designates the states $x,y$ that parametrises the forward and reverse statistics. This designation can be seen as a coarse-graining of the exact microstates of the system (that being the exact set of cards at hand), which is standard when defining states in physical experiments [8, 32, 10].
III.2.3 Work
Work has a straightforward meaning in the context of physics. Unfortunately, in the case of UNO, this is not as obvious. For this experiment, we designate the “work” $W$ as the number of deck interactions Alice makes from $x$ to $y$ ($y$ to $x$) for the forward (reverse) process. That is, whenever she needs to draw cards, $W$ increases by 1. Whenever she places a card, $W$ increases by 1 as well.
While other choices are definitely possible, this definition parallels the fact that work is fundamentally a measure of how long the experimenter acts on the target. Normally, the work parameter fluctuates because of the stochastic environment. Likewise, in our card-shedding context, Alice’s deck interactions fluctuates based on her decisions and the randomness coming from the other players and the deck. This definition of work, however, does not allow for negative work. Thus, in order to maintain Crooks’ notation and compare $P_{F}(W)$ with $P_{R}(-W)$ instead of $P_{R}(W)$, we define work in the reverse process to be negative.
III.2.4 Forward & Reverse Process
Finally, we need to define the forward and reverse processes. In a similar vein to other experiments with fluctuation relations, we characterize forward and reverse processes by the agent’s actions on the target for which the states are defined. For the context of this work, then, what defines the forward or reverse process has to do with how Alice makes decisions, changing the state of her hand. Hence, the forward process is designated as Alice playing according to a set of protocols that, statistically speaking, constitute a decent strategy at winning the game. For example, if Alice has a wildcard at hand, she will not play it unless she has no other legal moves. This helps her avoid the situation where she is forced to draw cards. The details of the protocol are given in Appendix A.
If this is the forward process, what then constitutes the reverse process? We designate it as the protocol that does the converse of the forward protocol. Corresponding to the example given before, Alice, under the reverse process, plays any wildcard she has on hand as soon as she can. As with the physical experiments, the work protocol (coming from the agent or experimentalist) is reversed while being done in the same stochastic environment (in this case being the other players).
III.3 Experimental Method
We perform our experiment by simulating $N=4\times 10^{6}$ UNO games on a computer and analysing the data of the results. Half of the games are played with normal UNO rules while in the other half the plus-cards (the +2 cards and +4 variation wildcards) are removed. Moreover, in half of the games Alice follows the forward algorithm, in the other half the reverse algorithm. In analysing a certain transition $x\rightarrow y$ we first obtain the probability distribution by sampling from the trajectories of Alice’s hand size in each game in the forward process games whenever this transition occurs, and vice versa for $y\rightarrow x$ in the reverse process games. We then compare the probability distributions $P_{f}(W|x\rightarrow y)$ and $P_{r}(-W|y\rightarrow x)$ to obtain the desired plots corresponding to the log of the ratio $P_{f}(W|x\rightarrow y)/P_{r}(-W|y\rightarrow x)$ against $W$. These graphs are shown in Fig. 3, one for ordinary UNO games and another for UNO games with the plus-cards removed. We see that Crooks’ relation is overall obeyed, with features to be analysed next.
IV Results
Fig. 3 shows the plots of $\ln{\left(\frac{P_{F}}{P_{R}}\right)}$ ) against $W$ from the collected statistics of the game for different transitions. There are two versions of the plot, one for the official UNO game, and another one for the same game but with the plus-cards removed. Let us first highlight the some notable features observed in the figures, which we will discuss in detail in the next section.
In all cases, it could be seen that a linear relation is obeyed in the $W\gtrsim 5$ regime, with a bend at around $W\sim 5$ which gives a hockey stick character to the plots. The behavior at larger values of $W$ seems to become more erratic, but this is due to the increasingly low number of samples. Regression lines for each of the transitions are also given in the graph. It can clearly be seen that the gradients are different for each of the transitions. Notably, the gradients can be negative. We will discuss the interpretation of this in Sec. V.2.
We now note differences between Fig.3(b) and Fig.3(a), which show the effects of removing the so-called plus-cards from the game. First, the gradients for the same $x,y$ transition decrease when such cards are removed from the game. Furthermore, for the ordinary deck, there are more data points due to the extension of the support to both even and odd $W$ values and there is a small oscillatory behaviour happening in the small $W$ regime. Finally, when the plus-cards are removed, there are fewer data points in the large $W$ regime. We elaborate more on these specific points in Sec. V.3.
V Discussion
Our numerical experiment has found Crooks’ relation occurs in a card game, for a non-thermodynamical analog of work and for dissipation in a finite non-Markovian bath. First, in order to gain some intuition on why Crooks’ theorem holds in such a case, we compare our process to the familiar Markovian random walk (Section V.1). We then trace the origin of the observed ‘temperature’ (be it positive or negative) to the relation between the starting point of the process and the typical hand size (Section V.2). Finally, we discuss how the plus-cards affect the extent in which Crooks’ relation is adhered to (Section V.3).
Figure 4: A comparison of $\ln{\left(\frac{P_{F}}{P_{R}}\right)}$ ) against $W$ before and after ’Markovianisation’. The particular transition in the figure is $5\rightarrow 2$. For other transitions, we observed qualitatively identical behaviors.
(a) UNO under official rules.
(b) UNO under official rules, without +2 cards and +4 variation wildcards.
Figure 5: The slopes $\beta(x\rightarrow y)$ of the linear regression lines for different transitions $x\rightarrow y$, for $1\leq x,y\leq 10$. In the original version of Crooks’ fluctuation theorem, this slope is the inverse temperature $\beta$ of the thermal bath.
Figure 6: Histogram of Alice’s hand size throughout all of the games under the complete official rules. Upper graph: forward process; lower graph: reverse process. Having sampled each process for over $N=10^{6}$ runs, the error bars are $\sqrt{p(1-p)/N}\approx 10^{-4}$, too small to be seen at the scale of the plot.
V.1 Comparison with a Markovian random walk
It is clear that, given our demarcations of the system, bath and designated state space, the process of playing UNO is non-Markovian. Indeed, the probability of losing/drawing a card on a given round depends not only on the number of cards at hand, but the actual cards themselves. Moreover, any given trajectory of an UNO game results in a memory that is embedded in the cards of the play deck, main deck and hands of other players. Transitions of the target system’s hand-size thus depend on this hidden memory, in a way that cannot be reduced to the previous, surface state, that it is transiting from.
Now, it may be insightful to quantify the extent of this non-Markovianity. To do this, we compare the original process to a ‘Markovianised’ version. The Markov transition matrix $M_{x,y}=P(y|x)$ will be given by the averaged transition probability of different hand sizes in a single round. We obtain these probabilities by averaging the probability of the $x\rightarrow y$ transition in a single round, regardless of the ‘actual’ state of the hand. What we obtain from this matrix is a Markov chain built from the UNO statistics where the probability of losing/drawing a card only depends on the current hand size. The Markov chain thus obtained is a finite, nonuniform random walk process. The nonuniformity can be understood through the fact that, in each round, Alice can only lose or draw one card, but the probabilities depend on the number of cards she currently has. For example, when she has 10 cards, it is more likely that she can play something compared to when she only has 2 cards. Thus, the transitions are dependent on position. It is finite since once she reaches state 0, the game ends.
With this, we rerun the simulations as before, but for the Markov chain. A comparison of the Crooks fluctuation graphs obtained before and after Markovianisation can be seen in Fig. 4. It is clear that the plots behave similarly in the small $W$ regime but differ as $W$ increases.
The analogy with a Markovian random walk may provide some intuition on the origin of the Crooks’ relation. Indeed, while a general treatment of a finite, nonuniform random walk is not trivial, it can be shown that Crooks’ theorem holds in an infinite, uniform random walk, when work corresponds to first passage time (see Appendix B).
With this, we see how the Crooks-like behavior emerges from the random-walk character in the intermediate $W$-regime. It is notable that this random-walk behaviour actually produces power law distributions in $W$ for each transition from $x$ to $y$ (3), as opposed to Gaussian distributions observed in physical fluctuation theorem experiments. This emphasizes that the settings which adhere to Crooks’ and Jarzynski’s theorems go beyond the typical thermodynamic case.
V.2 The analog of temperature
In the standard isothermal process, the slope of the Crooks line is given by the inverse temperature $\beta$. In our case, the “bath” is not a thermal bath, and this is most clearly manifested here: the slope of the Crooks line depends on the transition and may be negative: the values of $\beta(x\rightarrow y)$ are plotted in Fig. 5. For both versions of the game, we observe $\beta(x\rightarrow y)\approx-\beta(y\rightarrow x)$. For $x,y\gtrsim 4$, $\textrm{sgn}[\beta(x\rightarrow y)]=\textrm{sgn}[x-y]$: in words, the “temperature” is positive (negative) for the transitions that increase (decrease) Alice’s cards. The opposite behavior is observed when $x,y\lesssim 4$.
In order to understand this behavior, we plot the histogram of our coarse-grained variable (Alice’s hand’s size) throughout all the games, for both the forward and the reverse process (Fig. 6). We observe that the most probable values are 3, 4 and 5. There is also a slight difference between the forward and reverse processes, the maximum being 3 for the former and 4 for the latter.
With this, we can explain that the most obvious structure in Fig. 5 as follows: $\beta(x\rightarrow y)$ is positive for transitions that move the hand size away from the most probable number $\mu\approx 3\sim 4$, i.e. if $|y-\mu|>|x-\mu|$; it is negative for transition that bring the hand size closer to it, i.e. if $|y-\mu|<|x-\mu|$. Also, $|\beta|$ is typically larger if the transition happens far away from $\mu$.
V.3 The effects of plus-cards
As mentioned before, the removal of the plus-cards create some significant differences in the Crooks’ graph, which are: the extension of the support into both odd and even numbers, the oscillatory behaviour of the lines at small values of $W$, and the longer tails of the Crooks graphs. The first two are related and can be understood by our convention to count the action of drawing due to a +2 or +4 as a single deck interaction. Because of this, transitions that used to be possible in even (odd) steps can now be achieved in odd (even) steps too, given that Alice receives a plus-card. However, the probability of receiving a plus-card is relatively small, and most transitions still happen in the parity of $y-x$. This gives an oscillatory behaviour in both $P_{F}$ and $P_{R}$ between the two parity, and this behaviour carries over to the Crooks graphs. This behaviour, however, dies down at large values of $W$ as the chance of receiving a plus-card gets higher and higher the longer the trajectory is. Finally, the plus-cards also makes it harder to shed cards in general. Alice’s hand-size has a greater tendency to drift back and forth rather than decrease with regularity. This leads to generally longer paths and larger deck interactions.
Overall, the addition of plus-cards leads the game away from a nice smooth behaviour in the Crooks graphs. This is expected as the plus-cards give a jumping behaviour to otherwise smooth trajectories of the hand size.
We add that the same experiment was run for Crazy Eights (see Appendix C). Since it is played with standard playing cards without any analogue to plus-cards, the Crooks plots of Fig 7 have no significant oscillatory features, as expected.
VI Conclusion
We have shown that Crooks’ fluctuation relation can be found in non-physics scenarios, even in a game of cards. Using the card-shedding game of UNO, we first replaced the thermodynamical notion of work with the number of interactions of one player with the deck. The bath, into which heat is gained from and dissipated into, is replaced by the cards held by the other players or sitting in the deck. Then we defined the forward process as a behavior or protocol of a player that is relatively successful at shedding cards, and the reverse process as the converse of that forward protocol. With these definitions, we run simulations of the game and collected the statistics of “work” for both the forward and the reverse process. These statistics, which obey a power law, eventually show that the definition of work that we adopted obeys Crooks’ fluctuation theorem for those processes.
While a Crooks-like behavior is expected for a Markovian random walk, our process is non-Markovian, because the bath is finite and with memory. The unconventional nature of the bath is further demonstrated by the fact that the “temperature” (the inverse of the slope of the Crooks plot) can be either positive or negative, depending on the transition probabilities of the hand size. This behavior was explained as an artefact of the most common state being around the hand-size of $3$ and $4$.
Our work establishes another bridge connecting game theory and thermodynamics [33, 34, 35, 36], and provides a new example of the relevance of tools from stochastic thermodynamics outside thermal physics [37, 38]. Going forward, we notice that here we fixed a simple strategy for the player and focused on establishing an experiment in analogy with thermodynamical physics. It would be interesting to see a deeper connection between different strategies of the player and their effects on the fluctuation relation. Moreover, a more extensive survey of different games would be helpful in elucidating the significance and generality of Crooks’ fluctuation theorem.
The codes used to run the simulations can be found online  [^3] https://github.com/PeterSidajaya/uno-fluctuation.
Acknowledgments
We thank Alexia Auffèves, Francesco Buscemi, Andrés Ducuara, Hyukjoon Kwon, Dario Poletti and Antonios Varvitsiotis for helpful discussions and suggestions.
This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; and by the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002). This project started in the context of the Science Research Programme (SRP), a collaboration between the NUS Faculty of Science, NUS College of Design and Engineering, and the Gifted Education Branch of the Ministry of Education, Singapore. We also thank Chan Xin Yu for her support in this collaboration.
Appendix A Strategies used
In order to effect the forward and reverse processes, we created two algorithms, one that aims to win (forward) and another that aims to lose (reverse). Note that in UNO, it is typically legal to draw a card even when you can play a card from your hand. Thus, a possible reverse algorithm is just to draw a card every round. However, to make the reverse process less trivial we will force Alice to play when there is a possible legal move in her hand.
One of the strategy to win UNO is to keep any wildcard on the hand for as long as possible. This is because of the card’s flexibility to be played regardless of the state of the gamepile. Moreover, when playing a wildcard one should generally choose a colour that is most numerous in one’s hand. This is because colour changes is infrequent and colour tends to stay the same between subsequent rounds. These two observations constitute the strategy that Alice uses for the forward process. For the reverse strategy, Alice tries to throw away any wildcard she has as soon as possible. Moreover, she then chooses the least numerous colour on her hand, in the hope of being forced to draw a card as soon as possible.
The two algorithms are shown in Algorithm 1 and 2. Against three other bots playing completely randomly the win rate of the forward and reverse processes for the official UNO game are 33.8% and 24.7%, respectively.
if a card can be played then
       if there is a wildcard on hand then
             if a non-wildcard can be played then
                   play a random legal non-wildcard;
            else
                   $X\leftarrow$ most common colour on hand;
                   play the wildcard and declare $X$ as the next colour;
             end if
      else
             play a random legal card;
       end if
else
       draw a card $Y$;
       if $Y$ can be played then
             play $Y$;
       end if
end if
Algorithm 1 The forward process
if a card can be played then
       if there is a wildcard on hand then
            $X\leftarrow$ least common colour on hand;
             play the wildcard and declare $X$ as the next colour;
      else
             play a random legal card;
       end if
else
       draw a card $Y$;
       if $Y$ can be played then
             play $Y$;
       end if
end if
Algorithm 2 The reverse process
Appendix B Crooks’ relation for the time of first passage in an infinite random walk
Consider a random walk process on the lattice of integer numbers =$\mathds{Z}=\{...,-2,-1,0,1,2\}$. On each timestep, the probability of the walker going to the right ($+1$) is $p$, while the probability of it going to the left ($-1$) is $q=1-p$. In analogy with the definition of work adopted in the main text for the UNO game, let us define work as the first hitting/passage time from the initial to the final state, and denote this by the random variable $N$. Given an initial point $i$ in the lattice and given that the walker has moved from $i$ to $i+1$, what we want to find is the distribution of $P(N=2n-1|i\rightarrow i+1)$. By the method of generating function, it can be derived that [40]
$P(N=2n-1|i\rightarrow i+1)=\frac{(2n-3)!!}{n!}2^{n-1}p^{n}q^{n-1},$ (3)
where !!double-factorial!!!! is the double factorial, the product of all positive integers with the same parity as $n$.
So far this only concerns one process. To make a fluctuation relation, let us define two random walks on two different lattices, say $A$ and $B$, with different transition probabilities, say $p_{1},q_{1}$ and $p_{2},q_{2}$. Taking the logarithm of the ratio of the first passage times, we would quickly find out
$\ln\left(\frac{P_{A}(N=2n-1|i\rightarrow i+1)}{P_{B}(N=2n-1|i\rightarrow i+1)}\right)$
$=n\ln\left(\frac{p_{1}q_{1}}{p_{2}q_{2}}\right)-\ln\left(\frac{q_{1}}{q_{2}}\right)$
Taking analogies with Crooks, we can define
$\beta=\ln\left(\frac{p_{1}(1-p_{1})}{p_{2}(1-p_{2})}\right),\quad\Delta F=\frac{1}{\beta}\ln\left(\frac{(1-p_{1})}{(1-p_{2})}\right).$ (4)
As of this point, we have not yet elaborated on the relations of the two processes. To get a more meaningful interpretation, the second random walk should be the ‘reverse’ of the first one. For example, let $G(p,q)$ be the graph on which the forward process is done. Define $P_{F}=P(N|i\rightarrow i+1;G(p_{1},q_{1}))$ as the forward process. Next, define the reverse as $P_{R}=P(N|i+1\rightarrow i;G(p_{2},q_{2}))$. This is simply the first passage time of the walker going back to the initial state from the end state with a different set of transition probabilities. Note that
$P(N|i+1\rightarrow i;G(p_{2},q_{2}))=P(N|i\rightarrow i+1;G(q_{2},p_{2})).$
Now, if the forward and reverse processes are done on the same lattice, we would have $p_{2}:=p_{1}$ and $q_{2}:=q_{1}$. In this particular way of reversing, with Eq. (4) we will get $\beta=0$. In general, without context, there is no unique way to reverse a random walk process. However, as long as the reverse process is still a random walk, the linear Crooks-like relation of Eq. (4) will hold.
Appendix C Crazy Eights
(a) Crooks plot for Crazy Eights.
(b) Regression slopes for Crazy Eights.
Figure 7: The associated plots for the game Crazy Eights, of which UNO is based.
UNO is derived from a simpler game played with the standard playing card deck called the Crazy Eights. In the most basic version of the game, the eights act as the wildcards and the suites play the role of the colours. More homebrew rules can be added that make the game more similar to UNO. For completeness we create similar plots for the basic Crazy Eights game, which can be seen in Fig. 7. It can be seen that the plots are essentially the same, especially with UNO without the plus-cards (see Fig. 3(b)).
Appendix D A different kind of bath
Figure 8: The Crooks’ plot for an UNO game with other players playing smartly. Compare with Fig. 3(a).
In the main text, the other players other than Alice always play randomly without any strategy. This begs the question of whether this randomness of the other players ‘adds’ to the stochasticity to the bath and affects our result. The short answer to this question is: no. Fig 8 shows the Crooks’ plot obtained with the same method as in the main text but with all the other players using Algorithm 1 instead of playing randomly. As can be seen, there is little to no noticeable difference with Fig. 3(a).